# Continuous-Aperture Array (CAPA)-Based Wireless Communications: Capacity Characterization

### v1

@@ VII Conclusion
This article has proposed an analytically tractable framework for CAPA-based wireless communications by utilizing continuous operators to address the continuous nature of the EM field. This framework significantly differs from the conventional matrix-based transmission framework for SPDA-based communications. Based on this novel framework, we have characterized the capacity limits of uplink and downlink channels for CAPAs in both single-user and multiuser cases. For each investigated scenario, we derived closed-form expressions for the capacity-achieving detectors or source currents along with the corresponding capacity. Additionally, we have analyzed the channel capacity by specializing the derived results to several typical array structures. Through both theoretical analyses and numerical simulations, we have demonstrated that CAPA exhibits a higher sum-rate capacity and a more extensive capacity region compared to the conventional SPDA. This underscores CAPA’s potential as a promising transmission paradigm for future wireless networks.
@@
$\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}=\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}+{\mathsf{P}}_{2}\Big{\lvert}\int_{\mathcal{A}}\hat{\mathsf{H}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\Big{\rvert}^{2}.$
@@
×$\times{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}}{{(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{-\frac{3}{2}}\mathrm{d}x}\mathrm{d}z}.$
Subsequently, we can calculate the inner integral with the aid of [25, Eq. (2.264.5)] and the outer integral with the aid of [25, Eq. (2.284)], which yields the results in (60). The channel correlation factor can be written as follows:
$\rho=\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}{\mathsf{Q}_{1}^{*}(x,z)\mathsf{Q}_{2}(x,z)\mathrm{d}x}\mathrm{d}z}.$ (A44)
The above integrals can be calculated by applying the Chebyshev-Gauss quadrature rule, i.e., $\int_{-1}^{1}{\frac{f\left(x\right)}{\sqrt{1-x^{2}}}\mathrm{d}}x\approx\sum_{j=1}^{n}{f\left(x_{j}\right)}$ with $x_{j}=\cos\left(\frac{\left(2j-1\right)\pi}{2n}\right)$, lead to the results in (61). This completes the proof of Lemma 4.
@@
$\sum_{m_{x}\in\mathcal{M}_{x}}\sum_{m_{z}\in\mathcal{M}_{z}}\!\!\!{f_{\mathsf{sp}}^{(k)}\left(m_{x}\epsilon_{k},m_{z}\epsilon_{k}\right)\epsilon_{k}^{2}}\approx\!\!\iint_{{\mathcal{C}}_{k}}{f_{\mathsf{sp}}^{(k)}\left(x,z\right)\mathrm{d}x\mathrm{d}z}.$
@@
which can be calculated by utilizing [25, Eqs. (2.264.5) & (2.284)]. Furthermore, the expression of $\rho_{\mathsf{s}}$ follows from its definition, which completes the entire proof.

### v2

@@ VII Conclusion
This article has proposed an analytically tractable framework for CAPA-based multiuser communications by utilizing continuous operators to address the continuous nature of the EM field. This framework significantly differs from the conventional matrix-based transmission framework for SPDA-based communications. Based on this novel framework, we have characterized the capacity limits of uplink and downlink channels for CAPAs in both single-user and multiuser cases. For each scenario, we derived closed-form expressions for the capacity-achieving detectors or source currents along with the corresponding capacity. Additionally, we analyzed the channel capacity by specializing the derived results to several typical array structures. Through both theoretical analyses and numerical simulations, we have demonstrated that CAPA exhibits a higher sum-rate capacity and a more extensive capacity region compared to the conventional SPDA. This underscores CAPA’s potential as a promising transmission paradigm for future wireless networks.
Acknowledgement
The authors would like to thank the editor, Dr. Amine Mezghani, and anonymous reviewers for helpful comments that greatly improved the quality of the article.
@@ Acknowledgement
$\int_{\mathcal{A}}\lvert{\diamond}_{1}(\mathbf{x})\rvert^{2}{\rm{d}}{\mathbf{x}}=\int_{\mathcal{A}}\lvert{\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}})\rvert^{2}{\rm{d}}{\mathbf{r}}+{\mathsf{P}}_{2}\Big{\lvert}\int_{\mathcal{A}}\hat{\mathsf{H}}_{2}(\mathbf{r}){\mathsf{J}}_{{\mathsf{dl}},1}({\mathbf{r}}){\rm{d}}{\mathbf{r}}\Big{\rvert}^{2}.$ (A42)
@@
×$\times{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}}{{(x^{2}+z^{2}-2r_{k}\left(\Phi_{k}x+\Theta_{k}z\right)+r_{k}^{2})^{-\frac{3}{2}}\mathrm{d}x}\mathrm{d}z}.$ (A45)
Subsequently, we can calculate the inner integral with the aid of [56, Eq. (2.264.5)] and the outer integral with the aid of [56, Eq. (2.284)], which yields the results in (64). The channel correlation factor can be written as follows:
$\rho=\frac{1}{\sqrt{\mathsf{g}_{1}\mathsf{g}_{2}}}\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\int_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}{\mathsf{Q}_{1}^{*}(x,z)\mathsf{Q}_{2}(x,z)\mathrm{d}x}\mathrm{d}z}.$ (A46)
The above integrals can be calculated by applying the Chebyshev-Gauss quadrature rule, i.e., $\int_{-1}^{1}{\frac{f\left(x\right)}{\sqrt{1-x^{2}}}\mathrm{d}}x\approx\frac{\pi}{n}\sum_{j=1}^{n}{f\left(x_{j}\right)}$ with $x_{j}=\cos\left(\frac{\left(2j-1\right)\pi}{2n}\right)$, leading to the results in (65). This completes the proof of Lemma 4.
@@
$\sum_{m_{x}\in\mathcal{M}_{x}}\sum_{m_{z}\in\mathcal{M}_{z}}\!\!\!{f_{\mathsf{sp}}^{(k)}\left(m_{x}\epsilon_{k},m_{z}\epsilon_{k}\right)\epsilon_{k}^{2}}\approx\!\!\iint_{{\mathcal{C}}_{k}}{f_{\mathsf{sp}}^{(k)}\left(x,z\right)\mathrm{d}x\mathrm{d}z}.$ (A49)
@@
which can be calculated by utilizing [56, Eqs. (2.264.5) & (2.284)]. Furthermore, the expression of $\rho_{\mathsf{s}}$ follows from its definition, which completes the entire proof.